OpenStudy (anonymous):

Determine if the function is odd, even, or neither. h(x)=x^2-x^12 (^means raised. So x raised to the second, and x raised to the 12th) Thank you

OpenStudy (campbell_st):

so find f(-x) = (-x)^2 - (-x)^12 since both terms are even powers the result is f(-x) = x^2 - x^12 so f(-x) = f(x) which is the definition of an even function if f(-x) = -f(x) the function is odd... and if neither case occurs the function is neither odd nor even

OpenStudy (anonymous):

for a function to be odd: f(-x) = -f(x) so lets plug in (-x) h(-x) = (-x)^2 - (-x)^12 h(-x) = x^2 - x^12 = h(x) which means h(-x) does not equal -h(x) For a function to be even: f(-x) = f(x) h(-x) = (-x)^2 - (-x)^12 h(-x) = x^2 - x^12 h(-x) = h(x) yay! the function must be even.

OpenStudy (unklerhaukus):

even powers, y intercept 0 even function